# Constrained Online Convex Optimization with Polyak Feasibility Steps

### v1

@@ 4 Numerical Experiments
Although our primary contribution is our theoretical results, we also give numerical experiments to demonstrate the functionality of the algorithm and provide some empirical verification of the theoretical results. In these experiments, we benchmark the performance of our algorithm with the algorithm from Yu et al. (2017) as it has the best regret bound among those in Table 1.
We consider a $2$-dimensional toy setting with quadratic cost functions $f_{t}(\mathbf{x})=3\|\mathbf{x}-\mathbf{v}_{t}\|^{2}$ and linear constraints $A\mathbf{x}\leq\mathbf{b}$. We generate $\mathbf{v}_{t}$ by sampling uniformly from $[0,1]^{2}$, and take $A=[I\ -I]^{\top}$ and $\mathbf{b}=0.5\mathbf{1}$, where we use $I$ to denote the $2\times 2$ identify matrix. Therefore, we can define the constraint function as $g(\mathbf{x})=\max_{i\in[4]}\mathbf{a}_{i}^{\top}\mathbf{x}-b_{i}$, where $\mathbf{a}_{i}$ and $b_{i}$ are the $i$th row of $A$ and $\mathbf{b}$ respectively. [^5] Note that the constraint can also be written with the infinity-norm: $g(\mathbf{x})=\|x\|_{\infty}-b$. Accordingly, we take $G_{f}=\sqrt{2}$, $R=1$, $G_{g}=1$, $\epsilon=0.25$ and $\sigma=\frac{1}{\sqrt{2}}$.
@@
Figure 1: Simulation results for our algorithm (labeled PFS), alongside the algorithm from Yu et al. (2017) (labeled DPP). We also include a version of the algorithm in Yu et al. (2017) where the constraint is tightened (labeled DPP-T). The points indicate the average over 30 trials and the error bars and shading are ±$\pm 1$ standard deviation.
@@
Figure 2: Actions chosen by our algorithm and the one from Yu et al. (2017) in a simulation.
@@ 5 Conclusion
In this work, we give an algorithm for constrained OCO that uses Polyak feasibility steps to ensure anytime constraint satisfaction $g(\mathbf{x}_{t})\leq 0\ \forall t$ and $\mathcal{O}(\sqrt{T})$ regret, while only receiving feedback on the function value and subgradient at the played actions. We foresee this approach being relevant to safety-critical applications, where constraints must be satisfied despite having only limited constraint information.
@@ Appendix A Extension of Prior Work to No Cumulative Constraint Violation
In this section, we show how the prior work Yu et al. (2017) and Yuan and Lamperski (2018) can be extended to show cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ under Assumption 4 (which is the same as Assumption 1 in Mahdavi et al. (2012)). This follows a similar approach to Theorem 8 in Mahdavi et al. (2012), but we give it for completeness.
@@
Therefore, we have shown that the algorithm from Yu et al. (2017) can be extended to show $\mathcal{O}(\sqrt{T})$ regret and cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$, provided that Assumption 4 holds and that $T$ is sufficiently large, i.e. $T\geq\frac{C_{V}^{2}}{\epsilon^{2}}$. Note that the requirement that $T$ is sufficiently large is also required by other works that guarantee cumulative constraint satisfaction, e.g. Mahdavi et al. (2012); Jenatton et al. (2016).
We can use a similar process to give guarantees for Yuan and Lamperski (2018). The original guarantees in Yuan and Lamperski (2018) are of the form $\mathrm{Reg}_{T}\leq C_{R}T^{\max(\beta,1-\beta)}$ and $\sum_{t=1}^{T}[g(\mathbf{x}_{t})]_{+}\leq C_{V}T^{1-\beta/2}$. Therefore, choosing $\rho=\min\left(\epsilon,\frac{C_{V}}{T^{\beta/2}}\right)$ and taking $T\geq\frac{C_{V}^{2/\beta}}{\epsilon^{2/\beta}}$ ensures that $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$. Also, the resulting regret is,
@@
The order of the bound $\max(\beta,1-\beta,1-\beta/2)$ is minimized when $\beta=\frac{2}{3}$ and therefore, we get that $\mathrm{Reg}_{T}=\mathcal{O}(T^{2/3})$. Note that Yuan and Lamperski (2018) also gives a bound of the form $\sum_{t=1}^{T}([g(\mathbf{x}_{t})]_{+})^{2}\leq C_{V}T^{1-\beta}$, and therefore one might hope to guarantee $\sum_{t=1}^{T}([g(\mathbf{x}_{t})]_{+})^{2}\leq 0$. Unfortunately, the analysis approach does not immediately show this because the tightening parameter $\rho$ cannot be used to cancel out the violation as in (13).
Appendix B Missing proofs
B.1 Proof of Lemma 1
@@
Consider a closed convex function $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Assume that, for all $\mathbf{x}\in\mathbb{R}^{d}$, it holds that $\|\mathbf{s}\|\leq G$ for all $\mathbf{s}\in\partial h(\mathbf{x})$, and when $h(\mathbf{x})=0$, it holds that $\|\mathbf{s}\|\geq\sigma$ for all $\mathbf{s}\in\partial h(\mathbf{x})$. Then, take any $\mathbf{x}\in\mathbb{R}^{d}$ such that $h(\mathbf{x})>0$. Let $\mathbf{z}=\Pi_{\mathbf{x}:h(\mathbf{x})\leq 0}(\mathbf{x})$ and $\mathbf{s}_{x}\in\partial h(\mathbf{x})$. There exists $\gamma\geq 0$ and $\mathbf{s}_{z}\in\partial h(z)$ such that:
@@
$\sup_{\mathbf{v}\in\mathbb{B}}h(\mathbf{z}+\frac{\epsilon}{G}v)\leq h(\mathbf{z})+G\frac{\epsilon}{G}\leq-\epsilon+\epsilon=0.$
@@
Since $\mathbf{z}=\Pi_{\mathcal{S}}(\mathbf{x})$ and $\mathcal{S}$ is a closed convex set, we know that the vector $\mathbf{x}-\mathbf{z}$ is normal to $\mathcal{S}$. Furthermore, $\mathcal{S}=\{\mathbf{x}:h(\mathbf{x})\leq h(\mathbf{z})=0\}$ and the conditions of the lemma ensures that $\mathbf{0}$ is not a subgradient of $h$ at $\mathbf{z}$ (and therefore $\mathbf{z}$ is not a minimum of $h$) so Rockafellar (1970) (Corollary 23.7.1) tells us that there exists $\lambda\geq 0$ such that $\mathbf{x}-\mathbf{z}\in\lambda\partial h(\mathbf{z})$. Therefore, there exists $\gamma\geq 0$ and $\mathbf{s}_{z}\in\partial h(\mathbf{z})$ such that $\mathbf{x}-\mathbf{z}=\gamma\frac{\mathbf{s}_{z}}{\|\mathbf{s}_{z}\|}$.
@@
First note that when $g_{\rho}(\mathbf{x})\leq 0$, then $\mathbf{x}\in\mathcal{X}_{\rho}$ and therefore (8) holds with both sides zero. Next, we consider the case where $g_{\rho}(\mathbf{x})>0$. Due to the fact that $\mathbf{x}\in R\mathbb{B}$, and Assumptions 3 and 4, we can apply Lemma 3 #3 with $h\leftarrow g_{\rho}$, $G\leftarrow G_{g}$ and $\sigma\leftarrow\sigma$ to get that,
@@
B.2 Proof of (7)
In this section, we prove (7). This combines Fact 1 with the classical online gradient descent analysis from Zinkevich (2003).
@@ B.2 Proof of (7)
Given that $\mathbf{x}^{\star}\in\mathcal{X}\subseteq\mathcal{H}_{t}$ and $\mathcal{H}_{t}$ is convex, it holds that
$\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}$
$=\|\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t}))-\mathbf{x}^{\star}\|^{2}$
$\leq\|\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t})-\mathbf{x}^{\star}\|^{2}$
$\leq\|\mathbf{y}_{t}-\mathbf{x}^{\star}\|^{2}$
$\leq\|\mathbf{x}_{t}-\eta\nabla f_{t}(\mathbf{x}_{t})-\mathbf{x}^{\star}\|^{2}$
$=\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-2\eta\nabla f_{t}(\mathbf{x}_{t})^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})+\eta^{2}\|\nabla f_{t}(\mathbf{x}_{t})\|^{2}.$
@@
$\leq\frac{1}{2\eta}R^{2}+\frac{\eta}{2}G_{f}^{2}T$
where the third inequality uses Assumption 2. ∎

### v2

@@ 4 Numerical Experiments
Although our primary contribution is our theoretical results, we also give numerical experiments to demonstrate the functionality of the algorithm and provide some empirical verification of the theoretical results. [^6] Our experiment code is available at https://github.com/shutch1/OCO-Polyak-Feasibility-Steps. In these experiments, we benchmark the performance of our algorithm with the algorithm from Yu et al. (2017) as it has the best regret bound among those in Table 1.
We consider a $2$-dimensional toy setting with quadratic cost functions $f_{t}(\mathbf{x})=3\|\mathbf{x}-\mathbf{v}_{t}\|^{2}$ and linear constraints $A\mathbf{x}\leq\mathbf{b}$. We generate $\mathbf{v}_{t}$ by sampling uniformly from $[0,1]^{2}$, and take $A=[I\ -I]^{\top}$ and $\mathbf{b}=0.5\mathbf{1}$, where we use $I$ to denote the $2\times 2$ identity matrix. Therefore, we can define the constraint function as $g(\mathbf{x})=\max_{i\in[4]}\mathbf{a}_{i}^{\top}\mathbf{x}-b_{i}$, where $\mathbf{a}_{i}$ and $b_{i}$ are the $i$th row of $A$ and $\mathbf{b}$ respectively. [^7] Note that the constraint can also be written with the infinity-norm: $g(\mathbf{x})=\|x\|_{\infty}-b$. Accordingly, we take $G_{f}=\sqrt{2}$, $R=1$, $G_{g}=1$, $\epsilon=0.25$ and $\sigma=\frac{1}{\sqrt{2}}$.
Figure 2: Actions chosen by our algorithm and the one from Yu et al. (2017) in an experiment.
@@ 5 Conclusion
In this work, we give an algorithm for constrained OCO that uses Polyak feasibility steps to ensure anytime constraint satisfaction $g(\mathbf{x}_{t})\leq 0\ \forall t$ and $\mathcal{O}(\sqrt{T})$ regret, while only receiving feedback on the function value and subgradient at the played actions. We foresee this approach being particularly relevant to safety-critical applications, where constraints must be satisfied despite having only limited constraint information.
Acknowledgements
This work was supported by the National Science Foundation under grant #2330154.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Extension of Prior Work to No Cumulative Constraint Violation
In this section, we show how the prior work Yu et al. (2017) and Yuan & Lamperski (2018) can be extended to show cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ under Assumption 4 (which is the same as Assumption 1 in Mahdavi et al. (2012)). This follows a similar approach to Theorem 8 in (Mahdavi et al., 2012), but we give it for completeness.
@@
Therefore, we have shown that the algorithm from Yu et al. (2017) can be extended to show $\mathcal{O}(\sqrt{T})$ regret and cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$, provided that Assumption 4 holds and that $T$ is sufficiently large, i.e. $T\geq\frac{C_{V}^{2}}{\epsilon^{2}}$. Note that the requirement that $T$ is sufficiently large is also required by other works that guarantee cumulative constraint satisfaction, e.g. (Mahdavi et al., 2012; Jenatton et al., 2016).
We can use a similar process to give guarantees for Yuan & Lamperski (2018). The original guarantees in Yuan & Lamperski (2018) are of the form $\mathrm{Reg}_{T}\leq C_{R}T^{\max(\beta,1-\beta)}$ and $\sum_{t=1}^{T}[g(\mathbf{x}_{t})]_{+}\leq C_{V}T^{1-\beta/2}$. Therefore, choosing $\rho=\min\left(\epsilon,\frac{C_{V}}{T^{\beta/2}}\right)$ and taking $T\geq\frac{C_{V}^{2/\beta}}{\epsilon^{2/\beta}}$ ensures that $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$. Also, the resulting regret is,
@@
The order of the bound $\max(\beta,1-\beta,1-\beta/2)$ is minimized when $\beta=\frac{2}{3}$ and therefore, we get that $\mathrm{Reg}_{T}=\mathcal{O}(T^{2/3})$. Note that Yuan & Lamperski (2018) also gives a bound of the form $\sum_{t=1}^{T}([g(\mathbf{x}_{t})]_{+})^{2}\leq C_{V}T^{1-\beta}$, and therefore one might hope to guarantee $\sum_{t=1}^{T}([g(\mathbf{x}_{t})]_{+})^{2}\leq 0$. Unfortunately, the analysis approach does not immediately show this because the tightening parameter $\rho$ cannot be used to cancel out the violation as in (13).
Appendix B Comparison of Algorithm with Related Work
In this section, we discuss the differences between our algorithm design and the algorithm design in Zhang et al. (2025), which studies constrained variational inequalities. Although Zhang et al. (2025) studies a different setting and therefore uses a different analysis, there are some apparent similarities in algorithm design that we discuss here. To illustrate the similarities and differences, we can write our algorithm in a different form that allows for more direct comparison. In particular, when the cost functions are fixed $f_{t}=f$ and the constraint function is smooth, we can write our algorithm (with $\rho=0$) as,
$\mathbf{x}_{t+1}=\Pi_{R\mathbb{B}}\left(\mathbf{x}_{t}-\eta\nabla f(\mathbf{x}_{t})-\frac{[g(\mathbf{x}_{t})-\eta\nabla f(\mathbf{x}_{t})^{\top}\nabla g(\mathbf{x}_{t})]_{+}}{||\nabla g(\mathbf{x}_{t})||^{2}}\nabla g(\mathbf{x}_{t})\right),$ (14)
and the algorithm in Zhang et al. (2025) can be written as (see their Algorithm 2),
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\nabla f(\mathbf{x}_{t})-\eta\frac{[\alpha g(\mathbf{x}_{t})-\nabla f(\mathbf{x}_{t})^{\top}\nabla g(\mathbf{x}_{t})]_{+}}{||\nabla g(\mathbf{x}_{t})||^{2}}\nabla g(\mathbf{x}_{t}).$ (15)
In both algorithms, the step-size $\eta$ is $\Theta(\frac{1}{\sqrt{T}})$ and $\alpha=G_{g}/R$ in (15). As such, we can see several differences:
  • Zhang et al. (2025) applies the step-size $\eta$ to the entire third term in (15), while our algorithm in (14) applies the step size only to $\nabla f(\mathbf{x}_{t})^{\top}\nabla g(\mathbf{x}_{t})$ within the third term.
  • Zhang et al. (2025) applies the scaling $\alpha$ to $g(\mathbf{x}_{t})$ in (15), while our algorithm uses no such scaling in (14).
  • Zhang et al. (2025) uses an additional constraint to ensure that the iterates are bounded, whereas we use a projection on to $R\mathbb{B}$.
Appendix C Missing proofs
C.1 Proof of Remark 1
In this section, we show that Slater’s condition implies Assumption 4, provided that Assumption 1 holds.
Proposition 1.
Let Assumption 1 hold. Then, suppose that Slater’s condition holds, i.e. there exists positive real $\xi$ and $\mathbf{y}\in\mathbb{R}^{d}$ such that $g(\mathbf{y})\leq-\xi$. It follows that Assumption 4 holds with $\epsilon=c\xi$ and $\sigma=(1-c)\frac{\xi}{2R}$ for any $c\in(0,1)$.
Proof.
It holds that $\mathcal{X}^{\prime}=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})=-\epsilon\}$ is non-empty because there exists $\mathbf{z},\mathbf{y}$ such that $g(\mathbf{z})=0$ (given that $\mathcal{X}$ is nonempty and compact) and $g(\mathbf{y})\leq-\xi$. Specifically, since $-\epsilon\in[-\xi,0]$ and $g$ is continuous (via Corollary 10.1.1 in (Rockafellar, 1970)), it holds that there exists $\mathbf{z}$ such that $g(\mathbf{z})=-\epsilon$.
Then, we show that $\|\partial g(\mathbf{x})\|\geq\sigma$ for all $\mathbf{x}\in\mathcal{X}^{\prime}$. Indeed for all $\mathbf{x}\in\mathcal{X}^{\prime}$, it holds for all $\mathbf{s}\in\partial g(\mathbf{x})$ that,
$-\xi\geq g(\mathbf{y})\geq g(\mathbf{x})+\mathbf{s}^{\top}(\mathbf{y}-\mathbf{x})=-\epsilon+\mathbf{s}^{\top}(\mathbf{y}-\mathbf{x})\geq-\epsilon-\|\mathbf{s}\|\|\mathbf{y}-\mathbf{x}\|\geq-\epsilon-2\|\mathbf{s}\|R$
$\quad\implies\quad\|\mathbf{s}\|\geq\frac{\xi-\epsilon}{2R}=(1-c)\frac{\xi}{2R}=\sigma,$
where we use the fact that $\|\mathbf{y}-\mathbf{x}\|\leq 2R$ given that $\mathbf{x},\mathbf{y}\in\mathcal{X}\subseteq R\mathbb{B}$. ∎
C.2 Proof of Lemma 1
@@
Consider a closed convex function $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$, where we use the notation $\mathcal{S}=\{x\in\mathbb{R}^{d}:h(x)\leq 0\}$. Let $\mathcal{K}$ be a convex set such that $\mathcal{S}\subseteq\mathcal{K}$. Assume that, for all $\mathbf{x}\in\mathcal{K}$, it holds that $\|\mathbf{s}\|\leq G$ for all $\mathbf{s}\in\partial h(\mathbf{x})$, and when $h(\mathbf{x})=0$, it holds that $\|\mathbf{s}\|\geq\sigma$ for all $s\in\partial h(\mathbf{x})$. Then, take any $\mathbf{x}\in\mathcal{K}$ such that $h(\mathbf{x})>0$. Let $\mathbf{z}=\Pi_{\mathbf{x}:h(\mathbf{x})\leq 0}(\mathbf{x})$ and $\mathbf{s}_{x}\in\partial h(\mathbf{x})$. There exists $\gamma\geq 0$ and $\mathbf{s}_{z}\in\partial h(\mathbf{z})$ such that:
@@
$\sup_{\mathbf{v}\in\mathbb{B}}h(z+\frac{\epsilon}{G}\mathbf{v})\leq h(\mathbf{z})+G\frac{\epsilon}{G}\leq-\epsilon+\epsilon=0.$
@@
Since $\mathbf{z}=\Pi_{\mathcal{S}}(\mathbf{x})$ and $\mathcal{S}$ is a closed convex set, we know that the vector $\mathbf{x}-\mathbf{z}$ is normal to $\mathcal{S}$ at $\mathbf{z}$. Furthermore, $\mathcal{S}=\{\mathbf{x}:h(\mathbf{x})\leq h(\mathbf{z})=0\}$ and the conditions of the lemma ensures that $\mathbf{0}$ is not a subgradient of $h$ at $\mathbf{z}$ (and therefore $\mathbf{z}$ is not a minimum of $h$) so (Rockafellar, 1970) (Corollary 23.7.1) tells us that there exists $\lambda\geq 0$ such that $\mathbf{x}-\mathbf{z}\in\lambda\partial h(\mathbf{z})$. Therefore, there exists $\gamma\geq 0$ and $\mathbf{s}_{z}\in\partial h(\mathbf{z})$ such that $\mathbf{x}-\mathbf{z}=\gamma\frac{\mathbf{s}_{z}}{\|\mathbf{s}_{z}\|}$.
@@
We use the notation $g_{\rho}(x):=g(x)+\rho$. First note that when $g_{\rho}(\mathbf{x})\leq 0$, then $\mathbf{x}\in\mathcal{X}_{\rho}$ and therefore (8) holds with both sides zero. Next, we consider the case where $g_{\rho}(\mathbf{x})>0$. To do so, we will first show that the subgradient norm of $g_{\rho}(\mathbf{x})$ is lower bounded at all $\mathbf{x}$ such that $g_{\rho}(\mathbf{x})=0$. In particular, consider any $\mathbf{x}$ such that $g_{\rho}(\mathbf{x})=0$ (which necessarily satisfies $\mathbf{x}\in R\mathbb{B}$), and let $\mathbf{z}=\Pi_{g(\mathbf{x})\leq-\epsilon}(\mathbf{x})$ and $\mathbf{s}_{z}\in\partial g(\mathbf{z})$. Then, it follows from Assumption 4 that $\|\mathbf{s}_{z}\|\geq\sigma$ and therefore we can apply Lemma 3 #5 by setting $\mathcal{K}\leftarrow R\mathbb{B}$, $h\leftarrow g+\epsilon$, $\sigma\leftarrow\sigma$ to get that for all $\mathbf{s}_{x}\in\partial g(\mathbf{x})$,
$\|\mathbf{s}_{x}\|\geq\|\mathbf{s}_{z}\|\geq\sigma.$
Note that we have used the fact that $\partial(g(\mathbf{x})+\epsilon)=\partial g_{\rho}(\mathbf{x})=\partial g(\mathbf{x})$ since they only vary by a constant. Since we know that the subgradient norm of $g_{\rho}$ is lower bounded at the boundary of $\mathcal{X}_{\rho}$, we can apply Lemma 3 #3 with $\mathcal{K}\leftarrow R\mathbb{B}$, $h\leftarrow g_{\rho}$, $G\leftarrow G_{g}$ and $\sigma\leftarrow\sigma$ to get that,
@@
C.3 Proof of (7)
In this section, we prove (7). This combines Fact 1 with the classical online gradient descent analysis from (Zinkevich, 2003).
@@ C.3 Proof of (7)
Given that $\mathbf{x}^{\star}_{\rho}\in\mathcal{X}_{\rho}\subseteq\mathcal{H}_{t}$ and $\mathbf{x}^{\star}_{\rho}\in\mathcal{X}_{\rho}\subseteq\mathcal{X}\subseteq R\mathbb{B}$ (and both $\mathcal{H}_{t}$ and $R\mathbb{B}$ are convex), it holds that,
$\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}_{\rho}\|^{2}$ $=\|\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t+1}))-\mathbf{x}^{\star}_{\rho}\|^{2}$
$\leq\|\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t+1})-\mathbf{x}^{\star}_{\rho}\|^{2}$
$\leq\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}_{\rho}\|^{2}$
$\leq\|\mathbf{x}_{t}-\eta\nabla f_{t}(\mathbf{x}_{t})-\mathbf{x}^{\star}_{\rho}\|^{2}$
$=\|\mathbf{x}_{t}-\mathbf{x}^{\star}_{\rho}\|^{2}-2\eta\nabla f_{t}(\mathbf{x}_{t})^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star}_{\rho})+\eta^{2}\|\nabla f_{t}(\mathbf{x}_{t})\|^{2}.$
@@
$\leq\frac{2}{\eta}R^{2}+\frac{\eta}{2}G_{f}^{2}T$
where the third inequality uses Assumptions 1 and 2. ∎
C.4 Proof of Remark 2
In this section, we show that the guarantees of Corollary 1 hold under slightly weaker assumptions. In particular, we use the following Assumption 5 instead of Assumption 2. Assumption 5 is weaker than Assumption 2 in that it only requires that the cost gradients are bounded for all $\mathbf{x}\in\mathcal{X}$, rather than all $\mathbf{x}\in R\mathbb{B}$.
Assumption 5.
There exists positive real $G_{f}$ such that $\|\nabla f_{t}(\mathbf{x})\|\leq G_{f}$ for all $\mathbf{x}\in\mathcal{X}$ and $t\in[T]$.
We then state the guarantees as a proposition and prove it.
Proposition 2.
Let Assumptions 1, 3, 4 and 5 hold. Suppose that $g(\mathbf{x}_{1})\leq-\alpha$ for some $\alpha\in(0,\epsilon]$, and let $\rho=\frac{\alpha}{\sqrt{T}}$ and $\eta=\frac{\xi\rho}{G_{f}G_{g}}$, where $\xi=1-\sqrt{\gamma}$. It follows that the actions chosen by Algorithm 1 satisfy,
$\mathrm{Reg}_{T}$ $\leq\left(\frac{2G_{f}G_{g}R^{2}}{\xi\alpha}+\frac{G_{f}\xi\alpha}{2G_{g}}+\frac{G_{f}\alpha}{\sigma}\right)\sqrt{T},$
$g(\mathbf{x}_{t})$ $\leq 0\quad\forall t\in[T],$
Proof.
First, we show that $\|\nabla f_{t}(\mathbf{x}_{t})\|\leq G_{f}$ for all $t\in[T]$ by induction over the first $\tau$ rounds. The base case holds because $\mathbf{x}_{1}\in\mathcal{X}_{\rho}\subseteq\mathcal{X}$ and therefore $\|\nabla f_{t}(\mathbf{x}_{t})\|\leq G_{f}$ by Assumption 5. For the induction step, suppose that $\|\nabla f_{t}(\mathbf{x}_{t})\|\leq G_{f}$ for all $t\in[\tau]$. Then, note that the feasibility analysis in Section 3.4 only requires that the cost gradient is bounded at the previously played actions. In particular, this bound is applied in (11). Therefore, it holds that,
$g(\mathbf{x}_{\tau+1})\leq G_{g}\gamma^{t/2}\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})+\frac{\eta G_{g}G_{f}}{1-\sqrt{\gamma}}-\rho,$
implying that $g(\mathbf{x}_{\tau+1})\leq 0$ with the specified choice of algorithm parameters. Since $\mathbf{x}_{\tau+1}\in\mathcal{X}$, it holds that $\|\nabla f_{\tau+1}(\mathbf{x}_{\tau+1})\|\leq G_{f}$, and together with the induction hypothesis, it holds that $\|\nabla f_{t}(\mathbf{x}_{t})\|\leq G_{f}$ for all $t\in[\tau+1]$ and the induction step is complete. Thus, it holds that $\|\nabla f_{t}(\mathbf{x}_{t})\|\leq G_{f}$ for all $t\in[T]$. Along the way, we have also shown that $\mathbf{x}_{t}\in\mathcal{X}$ for all $t\in[T]$.
Next, we show that, because the cost gradients are bounded at the actions, the regret bound holds. In particular, the analysis of Term I in the regret analysis (see Section 3.3) holds because it only requires that $\|\nabla f_{t}(\mathbf{x}_{t})\|\leq G_{f}$ for all $t\in[T]$ (as shown in Section C.3). The analysis of Term II requires that $f_{t}(\Pi_{\mathcal{X}_{\rho}}(\mathbf{x}^{\star}))-f_{t}(\mathbf{x}^{\star})\leq G_{f}\|\Pi_{\mathcal{X}_{\rho}}(\mathbf{x}^{\star})-\mathbf{x}^{\star}\|$. This holds because the cost gradients are assumed to be bounded on $\mathcal{X}$ and therefore all $f_{t}$ are Lipschitz on $\mathcal{X}$. ∎